# High-Layer Attention Pruning with Rescaling

### v1

@@ Appendix B Baseline Method Pruning Layer Range Tuning
To show the superiority of our method compared to baseline methods, we utilize the perplexity metric to tune the pruning layer range for the baselines. For Mag-sp, Wanda-sp, and LLM-Pruner, the pruning ratio is the same across all pruning layers, whereas FLAP employs an adaptive pruning strategy, allowing different pruning ratios for each layer.
LLaMA3.1-8B [^6] https://huggingface.co/meta-llama/Llama-3.1-8B/blob/main/config.json and Mistral-7B-v0.3777https://huggingface.co/mistralai/Mistral-7B-v0.3/blob/main/config.json have 32 layers, each with 8 key/value attention heads. Qwen2-7B [^8] https://huggingface.co/Qwen/Qwen2-7B/blob/main/config.json consists of 28 layers with 4 key/value attention heads per layer, while Gemma2-9B [^9] https://huggingface.co/google/gemma-2-9b/blob/main/config.json includes 42 layers, each with 8 key/value attention heads. For LLaMA3.1-8B and Mistral-7B-v0.3, we allocate a KV cache budget of 3/4, while Qwen2-7B and Gemma2-9B have a KV cache budget of 6/7. This requires removing 64 key/value attention heads for LLaMA3.1-8B and Mistral-7B-v0.3, 16 for Qwen2-7B, and 48 for Gemma2-9B.
To determine the number of key/value attention heads to prune in each layer, we restrict the range of removable key/value attention heads to maintain integer pruning layer counts. For LLaMA3.1-8B and Mistral-7B-v0.3, 2 or 4 heads are removed per layer. Qwen2-7B allows 1 or 2 heads per layer, while Gemma2-9B permits 2, 3, 4, or 6 heads per layer. As a result, the pruning layer counts are configured as follows: for LLaMA3.1-8B and Mistral-7B-v0.3, the pruning layer number is 16 or 32; for Qwen2-7B, it is 8 or 16; and for Gemma2-9B, it is 8, 12, 16, or 24. For FLAP, where the pruning ratio can vary across layers, the pruning layer counts are configured as follows: for LLaMA3.1-8B and Mistral-7B-v0.3, the number of pruning layers is set to 12, 16, 20, 24, 28, or 32. For Qwen2-7B, the pruning layer counts are 8, 12, 16, 20, 24, or 28. Similarly, for Gemma2-9B, the options are 9, 12, 18, 24, 30, 36, or 42 layers.
For a fixed-length interval with undefined start and end points, we begin with the starting layer index set to 0. In each step, we increment both the starting layer index and the end layer index by 1 until the end layer index reaches the model’s highest layer. The results are presented in Table 6. Please note that we also integrate FLAP into HARP to prune FFN parameters. However, we do not tune the pruning layer range for our method.
@@
Appendix C Benchmark Details
For our evaluation, we employ the lm-evaluation-harness package (version 0.4.7) developed in [24]. It’s important to note that the lm-evaluation-harness provides two accuracy metrics: “acc” and “acc_norm” for the ARC-Challenge, OpenBookQA, and PIQA benchmarks. For these benchmarks, we report the “acc” accuracy results. Table 7 reports the number of tasks and the number of choices for each discriminative task. For the GSM8K and TriviaQA benchmarks, this package offers two accuracy metrics: “exact_match,strict-match” and “exact_match,flexible-extract”. In our reporting, we use the “exact_match,strict-match” accuracy results for these benchmarks. The number of tasks for GSM8K, NaturalQuestions, TriviaQA, MATH-hard, and BBH are 1319, 3610, 17944, 1324, and 6511, respectively. For long-context generation evaluation, we adopt the LongBench benchmark [7], which offers a comprehensive suite of tasks designed to assess the ability of language models to handle long sequences across diverse domains. LongBench includes 23 tasks categorized into six groups: (I) Single-Document QA, (II) Multi-Document QA, (III) Long In-Context Learning, (IV) Long Dialogue History Understanding, (V) Code Repository Understanding, and (VI) Long Structured Data Understanding. Each task varies in domain, document length, and complexity, providing a broad challenge spectrum. Expert performance ranges from $22\%$ to $89\%$ accuracy, with input lengths spanning from 13k to 167k tokens, and average human solving times between 5 and 13 minutes. This benchmark enables a rigorous evaluation of a model’s capacity to reason, retrieve, and generate over extended contexts.
@@ Appendix D Layer Selection with Heuristic Metric for Pruning Attention Heads
$\mathbf{I}(l)=\sum_{s\in\{q,k,v\}}\left(\sum_{i,j}\frac{\partial\mathbf{L}}{\partial\mathbf{W}^{s}_{ij}}\cdot\mathbf{W}^{s}_{ij}\right)^{2}$
@@
$\mathbf{S}_{A}^{(l)}=1-\mathbf{E}\left[\frac{\langle\mathbf{X}_{A}^{(l)},\mathbf{Y}_{A}^{(l)}\rangle}{\|\mathbf{X}_{A}^{(l)}\|_{2}\|\mathbf{Y}_{A}^{(l)}\|_{2}}\right]$
@@
$\mathbf{Y}_{A}^{l}=\mathbf{X}_{A}^{l}+\operatorname{Attention}\left(\operatorname{LayerNorm}\left(\mathbf{X}_{A}^{l}\right)\right).$
@@
Appendix E Limitations
One limitation of our approach is that a single unified $\alpha$ value does not generalize well across different models; instead, $\alpha$ needs to be tuned individually for each model. Additionally, we observe a significant performance drop on mathematical benchmarks, suggesting that further techniques are required to recover performance in these tasks. We leave this exploration to future work.

### v2

@@ Appendix B Baseline Method Pruning Layer Range Tuning
To fairly compare our method against existing pruning approaches, we use perplexity to tune the pruning layer range for all baseline methods. For Mag-sp, Wanda-sp, and LLM-Pruner, the pruning ratio is the same across all pruning layers, whereas FLAP employs an adaptive pruning strategy, allowing different pruning ratios for each layer.
LLaMA3.1-8B [^4] https://huggingface.co/meta-llama/Llama-3.1-8B/blob/main/config.json and Mistral-7B-v0.3555https://huggingface.co/mistralai/Mistral-7B-v0.3/blob/main/config.json have 32 layers, each with 8 key/value attention heads. Qwen2-7B [^6] https://huggingface.co/Qwen/Qwen2-7B/blob/main/config.json consists of 28 layers with 4 key/value attention heads per layer, while Gemma2-9B [^7] https://huggingface.co/google/gemma-2-9b/blob/main/config.json includes 42 layers, each with 8 key/value attention heads. For LLaMA3.1-8B and Mistral-7B-v0.3, we allocate a KV cache budget of 3/4, while Qwen2-7B and Gemma2-9B have a KV cache budget of 6/7. This requires removing 64 key/value attention heads for LLaMA3.1-8B and Mistral-7B-v0.3, 16 for Qwen2-7B, and 48 for Gemma2-9B.
To determine the number of key/value attention heads to prune in each layer, we restrict the range of removable key/value attention heads to maintain integer pruning layer counts. For LLaMA3.1-8B and Mistral-7B-v0.3, 2 or 4 heads are removed per layer. Qwen2-7B allows 1 or 2 heads per layer, while Gemma2-9B permits 2, 3, 4, or 6 heads per layer. As a result, the pruning layer counts are configured as follows: for LLaMA3.1-8B and Mistral-7B-v0.3, the pruning layer number is 16 or 32; for Qwen2-7B, it is 8 or 16; and for Gemma2-9B, it is 8, 12, 16, or 24. For FLAP, where the pruning ratio can vary across layers, we similarly tune the number of pruned layers to obtain strong perplexity under the same KV cache budget. Specifically, for LLaMA3.1-8B and Mistral-7B-v0.3, the number of pruning layers is set to 12, 16, 20, 24, 28, or 32. For Qwen2-7B, the pruning layer counts are 8, 12, 16, 20, 24, or 28. Similarly, for Gemma2-9B, the options are 9, 12, 18, 24, 30, 36, or 42 layers.
For a fixed-length interval with undefined start and end points, we begin with the starting layer index set to 0. In each step, we increment both the starting layer index and the end layer index by 1 until the end layer index reaches the model’s highest layer. The results are presented in Table 7. Note that we also integrate FLAP into HARP to prune FFN parameters. However, for our method we do not tune the pruning layer range and instead always prune the highest layers, which makes the comparison with tuned baselines conservative.
@@
Appendix C Benchmark Details
For our evaluation, we employ the lm-evaluation-harness package (version 0.4.7) developed in (Gao et al., 2024). This framework provides a unified interface to commonly used benchmarks and implements task-specific preprocessing and scoring. It is important to note that the lm-evaluation-harness provides two accuracy metrics, “acc” and “acc_norm”, for the ARC-Challenge, OpenBookQA, and PIQA benchmarks. Unless otherwise stated, for these benchmarks, we report the “acc” accuracy results. Table 8 reports the number of tasks and the number of choices for each discriminative task.
For the GSM8K and TriviaQA benchmarks, this package offers two accuracy metrics: “exact_match,strict-match” and “exact_match,flexible-extract”. In our reporting, we use the “exact_match,strict-match” accuracy results for these benchmarks. The number of tasks for GSM8K, NaturalQuestions, TriviaQA, MATH-hard, and BBH are 1319, 3610, 17944, 1324, and 6511, respectively.
For long-context generation evaluation, we adopt the LongBench benchmark (Bai et al., 2024), which offers a comprehensive suite of tasks designed to assess the ability of language models to handle long sequences across diverse domains. LongBench includes 23 tasks categorized into six groups: (I) Single-Document QA, (II) Multi-Document QA, (III) Long In-Context Learning, (IV) Long Dialogue History Understanding, (V) Code Repository Understanding, and (VI) Long Structured Data Understanding. Each task varies in domain, document length, and complexity, providing a broad challenge spectrum. According to Bai et al. (2024), expert performance ranges from $22\%$ to $89\%$ accuracy, with input lengths spanning from 13k to 167k tokens, and average human solving times between 5 and 13 minutes. This benchmark enables a rigorous evaluation of a model’s capacity to reason, retrieve, and generate over extended contexts.
@@ Appendix D Layer Selection with Heuristic Metric for Pruning Attention Heads
$\mathbf{I}(l)=\sum_{s\in\{q,k,v\}}\left(\sum_{i,j}\frac{\partial\mathbf{L}}{\partial\mathbf{W}^{s}_{ij}}\cdot\mathbf{W}^{s}_{ij}\right)^{2}.$
@@
$\mathbf{S}_{A}^{(l)}=1-\mathbf{E}\left[\frac{\langle\mathbf{X}_{A}^{(l)},\mathbf{Y}_{A}^{(l)}\rangle}{\|\mathbf{X}_{A}^{(l)}\|_{2}\|\mathbf{Y}_{A}^{(l)}\|_{2}}\right].$
@@
$\mathbf{Y}_{A}^{(l)}=\mathbf{X}_{A}^{(l)}+\operatorname{Attention}\left(\operatorname{LayerNorm}\left(\mathbf{X}_{A}^{(l)}\right)\right).$
@@
In our experiments, we use these heuristic metrics only for layer selection (i.e., to decide which layers are more suitable for pruning) and keep the head-level pruning and rescaling strategy fixed as described in the main text. This separation of concerns allows us to clearly attribute performance gains to the proposed high-layer attention pruning scheme rather than to a particular choice of heuristic.